OpenStudy (anonymous):

lisa deposits $6,900 at 2% for 1 year how much interest will she earn? what is the account balance at end of year?

OpenStudy (anonymous):

2% for the year?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

A = P(1 + .02) or $7038 = 6900(1+.02) with interest = 7038 - 6900 = $138